OpenStudy (anonymous):

Find the volume of a rectangular prism if the length is 3x2, the width is 2x, and the height is x2 + 3x + 5. Use the formula V = l ⋅ w ⋅ h, where l is length, w is width, and h is height, to find the volume. 5x5 + 18x4 + 30x3 5x6 + 18x4 + 30x3 6x5 + 18x4 + 30x3 6x6 + 18x4 + 30x3

OpenStudy (anonymous):

@MiniMato

OpenStudy (minimato):

\[V = l * w * h\] \[V = (3x^2) * (2x) * (x^2 + 3x + 5)\] First we can multiply the first 2 parts. \[(3x^2)(2x) = 6x^3\] Then we can multiply that by the last part \[(6x^3)(x^2 + 3x + 5)\] \[6x^5 + 18x^4 + 30x^3\]

OpenStudy (minimato):

So the answer would be C

OpenStudy (anonymous):

Oh wow it's that easy I was expecting something way out there

OpenStudy (anonymous):

Thanks @MiniMato

OpenStudy (anonymous):

Omg i jsut clicked the wrong one and that was the only question i missed

OpenStudy (minimato):

No problem. Problems like these may seem a bit hard at first when you look at it. But if you write it down on a sheet of paper or something, it becomes pretty easy if you know the steps :) You're welcome

OpenStudy (anonymous):

i picked b instead of c

OpenStudy (minimato):

Oh wow, sorry about that :( Well, at least you know how to do problems like those for next time

OpenStudy (anonymous):

Yep thanks though

OpenStudy (minimato):

Any time ^^ @Dante515